You hold a uniform, 28g pen horizontal with your thumb pushing down on one end and your index finger pushing upward 3.5cm from your thumb. The pen is 14 cm long.

-Find the two forces
Answer #1

as, pen is in static equlibrium so, net torque about index finger would be zero.

Torque = R x F

torque produce by thumb about index finger + torque produce by index fingure about itself + torque produce by mass of pen about index fingure = 0

taking clockwise positive direction and anticlockwise negative.

0.035 * F(thumb) + 0 = 0.07*Mg

F(thumb) = 0.54936 N

Net force also become zero as pen is in static equlibrium

F(Thumb) + F(Pen) =- F(index finger)

0.54936 N + 0.27468 N = - F(index finger)

F(index finger) = - 0.82404 N { opposite to force exert by thumb and mass of pen }